Please help me solve this problem: How many liters of pure alcohol must be added to 15 liters of a 20% solution to obtain a mixture which is 30% alcohol?
Answer by sofiyac(983)

let x be the liters of pure alcohol...pure alcohol is 100% alcohol so we have
20*15+100x=30(15+x) solve for x
300+100x=450+30x
70x=150 divide each side by 70
x=15/7 or 2.1 liters rounded